• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C. C. Chang

    Log entry time 00:09:47 on June27,2002

    Entry number 86848

    keyword=Shift Summary

    17:02 ROC14 hanged. Lost about 5 minutes of beam time.
    18:04 Finished Q2_k40; changed right arm to 58.42 deg and at 1.0745 GeV
    for Q2_klN.
    19:20 Ready for Q2_l50. Will first do the optics run, then reduce beam
    to 5 uA for harp run (with 9-fiol in place) before doing the
    efficiency and then final production run.
    19:50 Begin production run for Q2_l50.
    20:05 We forgot to run trigsetup before Q2_l50 run begins. So, the
    optics run (Run # 1833) and first (short) production run
    (Run # 1835) may not be good. The first real good run is
    Run # 1836. We will need to redo optics run at the end
    after doing efficiency run.
    20:28 MCC called to inform us that there was no beam due to
    vacuum problem. They don't know how long will it take to
    restore the beam.
    21:00 Called MCC to find out the beam status. They still have no
    idea as to when the beam will be back.
    22:05 MCC called to let us know that they have found the problem.
    Yet there is still no words as to when the beam will be back.
    23:50 Left arm Q1 tripped off.
    24:00 In the process of restoring left arm Q1. Still no sign of beam.